Assume that a country estimates its M1 money supply at $20
million. A broader measure of the money supply, M2, is $50 million.
The country’s gross domestic product is $100 million. Production or
real output for the country is 500,000 units or products.
a. Determine the velocity of money based on the M1 money
supply.
b. Determine the velocity of money based on the M2 money
supply.
c. Determine the average price for the real output.

Other checkable deposits $12 million
1
Expert's answer
2013-08-16T10:51:20-0400
a. We find the velosity from the Fisher's formula:
P*Q = M*V, GDP = M*V
V = M1/GDP = 20/100 = 0.2
b. V = M2/GDP = 50/100 = 0.5
c. P = GDP/Q = 100 million/500,000 = $200